OpenStudy (barrelracing):

write an equation in slope intercept form of the line that passes through the given point and is parallel to the graph of the given equation. (-4,-8) y = -2x+5

OpenStudy (diamondboy):

hi

OpenStudy (barrelracing):

hi

OpenStudy (diamondboy):

the line is parallel to the equation y=-2x+5 means they have d same slope

OpenStudy (diamondboy):

so wat do u think d slope wuld be?

OpenStudy (barrelracing):

well idk im not sure how to even start out

OpenStudy (diamondboy):

ok the equation is in d form y=mx+b where m is d slope.....so the value in front of x is the slope

OpenStudy (barrelracing):

so the slope is -2

OpenStudy (diamondboy):

yep correct!

OpenStudy (diamondboy):

now since d equation of the line we are looking for has that same slope we are done

OpenStudy (diamondboy):

well not done per say but almost there.......slope intercept form is y=mx+b

OpenStudy (barrelracing):

y= -2(-4) + 5(-8) y = 8 + -40

OpenStudy (diamondboy):

nope

OpenStudy (diamondboy):

y=-4 and x=-8 i.e -8= -2(-4) +b

OpenStudy (diamondboy):

now solve for b

OpenStudy (barrelracing):

im so confused

OpenStudy (diamondboy):

ok so where are u confused about

OpenStudy (barrelracing):

what am I suppose to plug into y=mx+b

OpenStudy (diamondboy):